### v1

# Purification of Noisy Measurements and Faithful Distillation of Entanglement

Abstract
We consider entanglement distillation in a realistic scenario with noisy operations in which quantum measurements that constitute a general quantum operation are particularly noisy. We present a protocol for purifying noisy measurements and show that with the help of the purification, imperfect local operations can be used to distill entanglement. We show that the purification protocol is robust against noise in implementation and analyze the purification in a practical realization: for measurement and gate errors up to $10\%$, we suggest that the purification with two additional qubits is cost-effective for distilling entanglement. The purification protocol is feasible with currently available quantum technologies and readily applied to entanglement applications.
Introduction. Entanglement is a resource that makes it possible to realize efficient information processing beyond the limitations of classical counterparts [1, 2]. Multipartite entangled states such as graph states can implement universal quantum computation by single-qubit measurements only [3, 4, 5]. Entanglement can establish secure communication over legitimate parties without computational assumptions [6, 7, 8, 9]. The capability of manipulating shared entangled states to useful ones by local operations of classical communication (LOCC) is a key to various entanglement applications in quantum information processing. In particular, maximally entangled states called entangled bits (ebits),
$|\phi^{+}\rangle^{(AB)}=\frac{1}{\sqrt{2}}(|00\rangle^{(AB)}+|11\rangle^{(AB)})$ (1)
are the unit of entanglement [10, 11, 12, 13, 14]. Sharing ebits allows parties to prepare other useful states, such as graph states, and to share entanglement at an arbitrary distance through quantum repeaters [15].
Distillation of entanglement is not only of practical interest to enable entanglement applications [16, 17, 18] but also of fundamental interest as it realizes an irreversible transformation from mixed or non-maximally entangled states to maximally entangled ones; note that the process is thus probabilistic [19, 20, 21, 22, 23]. Distillation of entanglement has been demonstrated experimentally [24, 25, 26, 27, 28, 29, 30, 31].
The currently available quantum technologies, the so-called noisy-intermediate-scale quantum (NISQ) technologies, contain noise in all steps, preparation, manipulation, and measurements [32]. In particular, errors are about ten times higher in a measurement than in a single-qubit gate [33, 34]. Hence, an LOCC protocol is also noisy. Note also that measurements are necessary in a construction of a general quantum operation, that can be realized by a unitary gate over a system and measurements on subsystems [35, 36].
Noisy LOCC may limit entanglement applications. For instance, the effect of noise on quantum operations has been analyzed in entanglement distillation protocols [37]. While the protocol contains two stages, bilateral controlled-NOT (CNOT) gates and measurements of single qubits, noise on gates or measurements prevents legitimate parties from distilling entanglement from weakly entangled states. The consequence reiterates the limitations of NISQ technologies towards entanglement applications such as the distribution of entangled states over a quantum network.
In this work, we present a purification protocol that probabilistically transforms a noisy measurement to a clean one, and apply it to entanglement distillation. We consider two protocols for distilling entanglement, one from unknown mixed states and the other from pure states, and show that in both cases, entanglement can be faithfully distilled with noisy measurements. The protocol is also robust against measurements and gate errors. We analyze minimal resources for practical applications of the purification protocol and show that two additional qubits suffice for distilling entanglement with noisy measurements.
Purifying a noisy measurement. We begin by presenting the protocol for purifying noisy measurements. The protocol shares some similarities with advantage distillation [38, 39, 40] in key agreement protocols. Throughout, we consider a qubit measurement in the computational basis $M_{i}=|i\rangle\langle i|$ for $i=0,1$. A noisy measurement may be described by a positive-operator-valued-measure (POVM), for $i=0,1$,
$\widetilde{M}_{i}=(1-p)|i\rangle\!\langle i|+p\frac{\mathds{1}}{2}$ (2)
where $p\in[0,1)$. With the decomposition $\mathds{1}=|0\rangle\langle 0|+|1\rangle\langle 1|$, one can see that $\widetilde{M}_{i}=(1-p/2)M_{i}+(p/2)M_{i\oplus 1}$, with $\oplus$ the bitwise addition, from which an error rate corresponds to $p/2$. For instance, an error rate in a NISQ system such as superconducting qubits is about $5\%$, meaning that a realistic parameter would be in a percent level, $p\approx 0.1$.
Let $\rho$ denote a qubit state on which a measurement is performed. A noiseless measurement shows probabilities $\mathrm{tr}[\rho M_{i}]$, and a noisy one finds $\mathrm{tr}[\rho\widetilde{M}_{i}]$ for $i=0,1$. We here devise a protocol to obtain error-free statistics $\mathrm{tr}[\rho M_{i}]$ with noisy measurements and additional qubits without improving experimental requirements. In effect, the protocol aims to transform a noisy POVM element $\widetilde{M}_{i}$ to a clean one $M_{i}$, even with some probability.
Figure 1: Detectors (red) contain noise in measurements, see Eq. (2). An error-suppressing detector (blue) that purifies a noisy measurement is devised by exploiting an interaction via a collective CNOT $V_{n}$ in Eq. (3) over $n$ qubits.
The protocol requires $(n-1)$ additional qubits and a collective CNOT gate $V_{n}$ as follows,
$V_{n}=|0\rangle\langle 0|\otimes\mathds{1}^{\otimes n-1}+|1\rangle\langle 1|\otimes X^{\otimes n-1}$ (3)
with a Pauli $X$ matrix. Then, $(n-1)$ additional qubits are prepared in a fiducial state $|0\rangle^{\otimes n-1}$ and a collective CNOT is applied,
$V_{n}~{}\rho\otimes|0\rangle\langle 0|^{\otimes n-1}~{}V_{n}^{\dagger}.$ (4)
Measurements on $n$ qubits are noisy. Suppose that $n$-bit outcomes are obtained, $x^{n}=(x_{1},x_{2},\cdots,x_{n})$ where $x_{k}\in\{0,1\}$ for $k=1,\cdots,n$. A POVM element describing the outcomes corresponds to
$\widetilde{M}_{x^{n}}=\widetilde{M}_{x_{1}}\otimes\cdots\otimes\widetilde{M}_{x_{n}}.$
Thus, we have the probability of outcome $x^{n}$ is found as follows,
$p(x^{n})=\mathrm{tr}[V_{n}~{}\rho\otimes|0\rangle\langle 0|^{\otimes n-1}~{}V_{n}^{\dagger}\widetilde{M}_{x^{n}}].$ (5)
Finally, we accept measurement results only when all outcomes are identical, i.e., either $0^{n}$ or $1^{n}$; otherwise, discard. Once outcomes are accepted, a probability of outcome $i$, for $i=0,1$, is given by,
$q^{(n)}(i)=\frac{p(i^{n})}{p(0^{n})+p(1^{n})}$ (6)
which converges to a probability with a noiseless measurement, i.e., $\mathrm{tr}[\rho M_{i}]$ as $n$ tends to be large.
To see that a probability in Eq. (6) can be arbitrarily close to a result of a noiseless measurement, we rewrite the probability of outcomes $x^{n}$,
$p(x^{n})$ $=$ $\mathrm{tr}[\rho\widetilde{Q}_{x_{1}}^{(n)}],~{}$ (7)
$\mathrm{where}$ $\widetilde{Q}_{x_{1}}^{(n)}=\langle 0_{2}\cdots 0_{n}|V_{n}^{\dagger}\widetilde{M}_{x^{n}}V_{n}|0_{2}\cdots 0_{n}\rangle$
for $x_{1}=0,1$. For instance, one can compute the following: $V_{n}|0_{2}\cdots 0_{n}\rangle=|0^{n}\rangle\langle 0|+|1^{n}\rangle\langle 1|$, and a POVM element on the first qubit when the outcomes are identical, $x^{n}\in\{0^{n},1^{n}\}$ is given by
$\widetilde{Q}_{i}^{(n)}$ $=$ $r_{0}^{(n)}M_{i}+r_{1}^{(n)}M_{i\oplus 1}$
$\mathrm{where}$ $r_{0}^{(n)}=(1-\frac{p}{2})^{n}~{}\mathrm{and}~{}r_{1}^{(n)}=(\frac{p}{2})^{n}$ (8)
for $i=0,1$. From Eq. (6), POVM elements on the first qubit when all outcomes are identical are given by,
$q^{(n)}(i)$ $=$ $\mathrm{tr}[\rho Q_{i}^{(n)}],~{}\mathrm{where}~{}$ (9)
$Q_{i}^{(n)}$ $=$ $\frac{(1-\frac{p}{2})^{n}}{p(0^{n})+p(1^{n})}M_{i}+\frac{(\frac{p}{2})^{n}}{p(0^{n})+p(1^{n})}M_{i\oplus 1}$
Since $p(0^{n})+p(1^{n})=(1-\frac{p}{2})^{n}+(\frac{p}{2})^{n}$, we have achieved a POVM element $Q_{i}^{(n)}$ which arbitrarily converges to a clean one $M_{i}$ as $n$ tends to be large. Equivalently, the fidelity between $Q_{i}^{(n)}$ and $M_{i}$ can be arbitrarily close to $1$ as $n$ increases, i.e.,
$\mathrm{tr}[Q_{i}^{(n)}M_{i}]\rightarrow 1~{}~{}\mathrm{as}~{}~{}n\rightarrow\infty$ (10)
Hence, we have shown a noisy measurement can be purified. It is also worth mentioning that the purification works for $p\in(0,1)$.
Distilling entanglement from mixed states. We now consider the distillation of entanglement with noisy local operations. We emphasize that entanglement distillation is key to various quantum information applications, such as entanglement distribution over a quantum network and measurement-based quantum computing. In what follows, we show how the purification of a noisy measurement enables existing distillation protocols to distill entanglement from all entangled states, even if local operations constructed by noisy measurements are imperfect.
Let us briefly summarize the two-way distillation protocol. For copies of states shared by two parties, called Alice and Bob, the protocol begins by twirling shared and unknown states so that two parties share an isotropic state in a single-copy level, characterized as follows,
$\rho_{F}=F|\phi^{+}\rangle\langle\phi^{+}|+\frac{1-F}{3}(\mathds{1}-|\phi^{+}\rangle\langle\phi^{+}|),$ (11)
where $F=\langle\phi^{+}|\rho_{F}|\phi^{+}\rangle$ is called a singlet fraction. Then, taking two copies $\rho_{F}^{A_{1}B_{1}}\otimes\rho_{F}^{A_{2}B_{2}}$, two parties apply bilateral CNOT gates, $V_{2}^{A_{1}A_{2}}$ and $V_{2}^{B_{1}B_{2}}$, see Eq. (4), and perform measurements in the second register $A_{2}B_{2}$ in the computational basis $\{M_{0},M_{1}\}$. If outcomes are equal, a state remaining in the first register $A_{1}B_{1}$ is accepted. The singlet fidelity of the accepted state is given by
$F^{\prime}=\frac{F^{2}+\left(\frac{1-F}{3}\right)^{2}}{F^{2}+2F\left(\frac{1-F}{3}\right)+5\left(\frac{1-F}{3}\right)^{2}}.$ (12)
We have $F^{\prime}>F$ for $F>1/2$, which is also the condition that a state $\rho_{F}$ is entangled. Therefore, the protocol above can distill entanglement from all two-qubit entangled states; hence, for two-qubit states, entanglement implies distillability.
Figure 2: Two parties Alice and Bob apply error-suppressing detectors for distilling entanglement. For a realistic measurement error rate up to $10\%$, two additional qubits on each party immediately allow entanglement distillation from weakly entangled states up to a singlet fraction $F>0.503$.
When a measurement in the second register $A_{2}B_{2}$ is noisy, the local operations of two parties are imperfect. Consequently, entanglement cannot be distilled from weakly entangled states. For instance, for a noisy measurement with $p=0.1$, see Eq. (2), entanglement can be distilled only when $F>0.617$ [37]. No entanglement can be distilled if $p>1-1/\sqrt{2}\approx 0.29$.
We first apply the purification of a noisy measurement to enable entanglement distillation and then show the robustness of the protocol by considering noise that may be contained in CNOT gates. The purification protocol will attain the entanglement distillability from two-qubit states. For a pair of isotropic states obtained after twirling shared states, two parties perform bilateral CNOT gates followed by noisy measurements. Alice and Bob exploit $n-1$ and $m-1$ additional qubits, respectively, for purifying noisy measurements. A resulting singlet fidelity of Alice and Bob relies on measurements on $n$ ad $m$ qubits and is given by,
$F^{\prime}=\frac{F^{2}+\left(\frac{1-F}{3}\right)^{2}+g^{(n,m)}(F)}{F^{2}+2F\left(\frac{1-F}{3}\right)+5\left(\frac{1-F}{3}\right)^{2}+4g^{(n,m)}(F)}$
where $g^{(n,m)}(F)$ is defined as follows,
$g^{(n,m)}(F)$ $=$ $\left(\frac{F(1-F)}{3}+\left(\frac{1-F}{3}\right)^{2}\right)\frac{r_{odd}^{(n,m)}}{r_{even}^{(n,m)}},$ (13)
$r_{even}^{(n,m)}$ $=$ $(1-\frac{p}{2})^{n}(1-\frac{p}{2})^{m}+(\frac{p}{2})^{n}(\frac{p}{2})^{m},~{}\mathrm{and}$
$r_{odd}^{(n,m)}$ $=$ $(1-\frac{p}{2})^{n}(\frac{p}{2})^{m}+(\frac{p}{2})^{n}(1-\frac{p}{2})^{m}.$ (14)
As $n$ and $m$ tends to be large, we have that $g^{(n,m)}(F)$ converges to zero and reproduce the fidelity in Eq. (12) of a noiseless instance.
Figure 3: In a noiseless case, two-qubit states having a singlet fraction $F>1/2$ are distillable. With noisy measurements with a noise fraction $p=0.1$, the distillation protocol works for entangled states with a singlet fraction $F>0.617$. The purification of noisy measurements leads to an improvement: entanglement can be distilled from entangled states with $F>0.505$. The protocol is also robust: even for measurements having different error rates $p_{i}\in(0.025,0.175)$, see Eq. (16), the protocol faithfully increases a singlet fidelity (grey).
For the full generality in error rates on the measurements, we suppose that error rates in independent individual measurements are not identical. Let $p_{k}^{A}$ denote a noise fraction of the $k$-the measurement of Alice and $p_{k}^{B}$ for Bob, see Eq. (2). The resulting fidelity is obtained with $g^{(n,m)}(F)$ in Eq. (13) defined by functions $r_{even}^{(n,m)}$ and $r_{odd}^{(n,m)}$,
$r_{even}^{(n,m)}$ $=$ $\prod_{i=1}^{n}\left(1-\frac{p^{A}_{i}}{2}\right)\prod_{i=1}^{m}\left(1-\frac{p^{B}_{i}}{2}\right)+$ (15)
$\prod_{i=1}^{n}\left(\frac{p^{A}_{i}}{2}\right)\prod_{i=1}^{m}\left(\frac{p^{B}_{i}}{2}\right),~{}\mathrm{and}$
$r_{odd}^{(n,m)}$ $=$ $\prod_{i=1}^{n}\left(1-\frac{p^{A}_{i}}{2}\right)\prod_{i=1}^{m}\left(\frac{p^{B}_{i}}{2}\right)+$ (16)
$\prod_{i=1}^{n}\left(\frac{p^{A}_{i}}{2}\right)\prod_{i=1}^{m}\left(1-\frac{p^{B}_{i}}{2}\right).$
We detail the derivation in Appendix A. In Fig. 3, we consider an instance with $p_{i}\in(0.025,0.175)$ and show that entanglement increases faithfully.
Then, the distillation protocol increases entanglement whenever $F^{\prime}>F$, for which we compute:
$F^{\prime}-F$ $=$ $-8\left(F-\frac{1}{4}\right)\left(F-L^{(n,m)}\right)\left(F-1\right)$
$\mathrm{where}$ $L^{(n,m)}=\frac{r_{even}^{(n,m)}+r_{odd}^{(n,m)}}{2\left(r_{even}^{(n,m)}-r_{odd}^{(n,m)}\right)}.$ (17)
Thus, we have $F^{\prime}>F$ for $F\in(L^{(n,m)},1)$ where the lower bound $L^{(n,m)}$ depends on the numbers of measurement outcomes $n$ and $m$ for purifying measurements and a noise fraction $p$ but not an initial fidelity $F$. Note that $L^{(n,m)}$ converges to $1/2$ as $n$ and $m$ tend to be large; hence, the noiseless instance is reproduced.
Examples. We consider entanglement distillation with noisy measurements in a realistic scenario. We are interested in the lower bound $L^{(n,m)}$ to see how the purification can push the limitations in distillability.
For noisy measurements with $p=0.2$, i.e., an error rate $10\%$, we have lower bounds as follows, n 1 2 3 4 $L^{(n,1)}$ 0.781 0.640 0.627 0.625 $L^{(n,2)}$ 0.640 0.525 0.514 0.513 $L^{(n,3)}$ 0.627 0.514 0.503 0.502 From $L^{(n,1)}$, we conclude the purification on noisy measurements of a single party yet achieves a limited range: it improves up to $0.625$ by the purification of noisy measurements of Alice. We also find that from $L^{(n,2)}$ and $L^{(n,3)}$, two additional qubits, i.e., $(n,m)=(3,3)$, are cost-effective as the lower bound is immediately reach $0.503$.
Then, we consider measurements with a noise fraction $p=0.1$, meaning that an error rate is $5\%$, we have lower bounds for distillability as follows, n 1 2 3 4 $L^{(n,1)}$ 0.617 0.559 0.556 0.556 $L^{(n,2)}$ 0.559 0.505 0.500 0.500 which shows that the purification of noisy measurements with single additional qubits on two parties nearly recovers the noiseless bound $1/2$, see Fig. 3 in which the improvement of lower bounds is shown.
Robustness against noise. For the robustness of the purification protocol, we consider noise in a collective CNOT gate $V_{n}$, which may be decomposed as
$V_{n}^{(S_{1}\cdots S_{n})}=\bigotimes_{j=2}^{n}V_{2}^{(S_{1}S_{j})}.$ (18)
Since a quantum operation is implemented by single-qubit and CNOT operations in practice, we consider a noise fraction $\epsilon$ in a CNOT gate [41],
$(1-\epsilon)V_{2}(~{}\cdot~{})V_{2}^{\dagger}+\epsilon\frac{\mathds{1}}{2}\otimes\frac{\mathds{1}}{2}.$ (19)
Then, a measurement is described by $r_{i}^{(n)}$ for $i=0,1$ in Eq. (8) that can be expressed by a recurrence relation,
$r_{0}^{(n+1)}$ $=(1-\epsilon)r_{0}^{(n)}(1-\frac{p}{2})+\frac{\epsilon}{4}(r_{0}^{(n)}+r_{1}^{(n)})~{}\mathrm{and}$
$r_{1}^{(n+1)}$ $=(1-\epsilon)r_{1}^{(n)}(\frac{p}{2})+\frac{\epsilon}{4}(r_{0}^{(n)}+r_{1}^{(n)}).$
Since errors of a measurement and a CNOT gate are comparable, we consider noise fractions $p=0.1$ in measurements and $\epsilon=0.1$ in the purification protocol and find that one can reach a lower bound $0.566$ with two additional qubits as follows, n 1 2 3 4 $L_{\epsilon}^{(n,n)}$ 0.617 0.570 0.566 0.566 where $L_{\epsilon}^{(n,n)}$ denotes the lower bound of an initial fidelity over which entanglement can be distilled. An analytic expression on the lower bound is shown in Appendix B. We emphasize that two additional qubits in the purification of noisy measurements are sufficient for a practical purpose.
Distilling entanglement from pure states. We next consider cases where two parties share pure states which are not maximally entangled. From the Schmidt decomposition, let us suppose that two parties share copies of a state in the following,
$|\psi(\theta)\rangle^{(AB)}=\sin\theta|00\rangle^{(AB)}+\cos\theta|11\rangle^{(AB)}$
where $\theta\in(0,\frac{\pi}{4})$ so that $\cos\theta>\sin\theta$. In a noiseless scenario, Bob implements a local operation with Kraus operators,
$K_{0}=|0\rangle\!\langle 0|+\tan\theta|1\rangle\!\langle 1|~{}\mathrm{and}~{}K_{1}=\sqrt{1-\tan^{2}\theta}|1\rangle\!\langle 1|.~{}~{}~{}~{}~{}~{}~{}$ (20)
Kraus operators above are realized with the help of an additional system denoted by $E$. Let us introduce a controlled-$W$ gate on systems $BE$
$U^{(BE)}$ $=$ $|0\rangle\!\langle 0|^{(B)}\otimes I^{(E)}+|1\rangle\!\langle 1|^{(B)}\otimes W^{(E)},~{}\mathrm{where}$
$W$ $=$ $\begin{pmatrix}\tan\theta&-\sqrt{1-\tan^{2}\theta}\\ \sqrt{1-\tan^{2}\theta}&\tan\theta\end{pmatrix},$ (21)
such that $W|0\rangle=\tan\theta|0\rangle+\sqrt{1-\tan^{2}\theta}|1\rangle$. For a state $\rho^{(BE)}$, a measurement is performed on system $E$ after a controlled-$W$ gate. An outcome $m\in\{0,1\}$ realizes a Kraus operator $K_{m}$ in Eq. (20) on system $B$, i.e.,
$K_{m}^{(B)}=_{E}\!\!\langle m|U^{(BE)}|0\rangle_{E}.$ (22)
When an outcome 00 occurs and thus $K_{0}$ is realized, an ebit is distilled. Otherwise, two parties share a product state since the other Kraus operator $K_{1}$ is rank-one. The probability of having outcome 00 is given by $2\sin^{2}\theta$, i.e., we have
$(\mathds{1}\otimes K_{0})|\psi(\theta)\rangle=\sqrt{2}\sin\theta|\phi^{+}\rangle.$ (23)
Thus, entanglement can be distilled by local operations with some probability, for which a measurement on a subsystem is essential.
Figure 4: Two parties sharing a pure state can distill an ebit by applying a local filtering operation on one of the parties, where the local operation is composed of a controlled-$W$ gate in Eq. (21) followed by an error-suppressing detector that purifies a noisy measurement.
If a measurement is noisy, Kraus operators in Eq. (22) cannot be precisely realized and consequently entanglement is not distilled. In particular, two parties may share mixed states due to noise existing in Kraus operators. Similarly to Eq. (23), we have that
$\mathrm{tr}_{E}[U^{(BE)}|\psi\rangle\langle\psi|^{(AB)}\otimes|0\rangle\langle 0|^{(E)}U^{(BE)\dagger}\widetilde{M}_{0}^{(E)}]=$ (24)
$(1-\frac{p}{2})K_{0}^{(B)}|\psi\rangle\langle\psi|^{(AB)}K_{0}^{(B)\dagger}+\frac{p}{2}K_{1}^{(B)}|\psi\rangle\langle\psi|^{(AB)}K_{1}^{(B)\dagger}$
where a measurement $\widetilde{M}_{0}^{(E)}$ is noisy, see Eq. (2). Thus, a Kraus operator in Eq. (22) when a measurement is noiseless no longer applies. Even if a measurement outcome 00 occurs, the other rank-one Kraus operator $K_{1}$ appears with a probability $p/2$.
Purification of a noisy measurement applies to $\widetilde{M}_{0}$ in Eq. (24) with $n-1$ additional qubits, and let $\rho_{n}^{(AB)}$ denote the resulting state after Bob’s operation when $n$ measurement outcomes are identical. Once a resulting state is accepted, the fidelity is given by,
$F_{n}$ $=$ $\langle\phi^{+}|\rho_{n}^{(AB)}|\phi^{+}\rangle$
$=$ $\frac{1}{p_{succ}}\left({(1-\frac{p}{2})^{n}2\sin^{2}\theta+\frac{1}{2}(\frac{p}{2})^{n}(1-2\sin^{2}\theta)}\right)$
where $p_{succ}$ is the probability of $n$ identical outcomes in the purification protocol,
$p_{succ}=(1-\frac{p}{2})^{n}2\sin^{2}\theta+(\frac{p}{2})^{n}(1-2\sin^{2}\theta).$
As $n$ tends to be large, the fidelity converges to $1$, i.e.,
$F_{n}\rightarrow 1~{}~{}\mathrm{as}~{}~{}n\rightarrow\infty.$
Thus, we have shown distillation of an ebit from a pure state by purifying a noisy measurement. In fact, $F_{n}$ converges to $1$ as $n$ increase, e.g., $p=0.1$ and $\theta=\pi/16$: n 1 2 3 4 $F_{n}$ 0.805 0.984 0.999 1.000 If noise is present in an implementation of a collective CNOT gate, see Eqs. (18) and (19), one cannot reach a unit fidelity. For instance, for a noise fraction $p=0.1$ and shared state $\theta=\pi/16$ we have n 1 2 3 4 $F_{n}$ 0.805 0.914 0.924 0.924 showing that a fidelity converges to $0.924$ with two additional qubits. We detail analytic derivations in Appendix C.
Conclusion. We have shown that a noisy measurement can be purified and applied to the distillation of entanglement. For both cases of mixed and pure states, we have shown that entanglement can be distilled by purifying noisy measurements. The purification protocol is robust against measurement and gate errors in implementation. For a realistic scenario where an error rate is about $5$-$10\%$, we have shown that two additional qubits for the purification suffice almost to recover the noiseless range for distillability: the single fraction of $F>1/2$. Our results are a versatile tool to suppress measurement errors for realizing practical entanglement applications with currently available quantum technologies.
In future investigations, it would be interesting to consider noise in bilateral CNOT gates in the distillation of entanglement. Noise in bilateral CNOT gates prevents two parties from sharing an ebit: even a tiny fraction of noise disallows two parties to establish an ebit. We note that the proposed purification protocol is yet of no use in dealing with errors from bilateral CNOT gates.
This work is supported by the National Research Foundation of Korea (Grant No. NRF-2021R1A2C2006309, NRF-2022M1A3C2069728), the Institute for Information & Communication Technology Promotion (IITP) (RS-2023-00229524), and grant-in-aid of LG Electronics.
Appendix A Derivation of Eq. (16)
In this section, we take different error rates in noisy measurements into account by generalizing Eq.(8), where all measurements have an identical error rate, to cases where each measurement has a different error probability. Let $p_{i}$ denote a noise fraction of the $i$-th measurement, so that
$\widetilde{M}_{x^{n}}=\widetilde{M}_{x_{1}}\otimes\cdots\otimes\widetilde{M}_{x_{n}},~{}\text{ where }~{}\widetilde{M}_{x_{i}}$ $=(1-\frac{p_{i}}{2})M_{x_{i}}+(\frac{p_{i}}{2})M_{x_{i}\oplus 1}.$ (25)
Recall that a POVM element is found as,
$\widetilde{Q}_{x}^{(n)}$ $=\langle 0_{2}\cdots 0_{n}|V_{n}^{\dagger}\widetilde{M}_{x^{n}}V_{n}|0_{2}\cdots 0_{n}\rangle=r_{0}^{(n)}M_{x}+r_{1}^{(n)}M_{x\oplus 1}.$
In the above, coefficients $r_{0}^{(n)}$ and $r_{1}^{(n)}$ can be obtained recursively from an initial condition $\widetilde{Q}_{x}^{(1)}=\widetilde{M}_{x_{1}}$, where $r_{0}^{(1)}=(1-\frac{p_{1}}{2})$ and $r_{0}^{(1)}=\frac{p_{1}}{2}$, and updating $n$. That is, we exploit the following relation in the update $k=2,\cdots,n$:
$\widetilde{Q}_{x}^{(k)}$ $=\langle 0_{k}|V_{2}^{\dagger}\widetilde{Q}_{x}^{(k-1)}\otimes\widetilde{M}_{x_{k}}V_{2}|0_{k}\rangle,$ (26)
where $V_{2}$ is a CNOT gate that takes the first qubit as a control and the $k$-th qubit as a target. From the equations (25, 26), we have
$\widetilde{Q}_{x}^{(k)}$ $=\prod_{i=1}^{k}\left(1-\frac{p_{i}}{2}\right)|0\rangle\!\langle 0|+\prod_{i=1}^{k}\left(\frac{p_{i}}{2}\right)|1\rangle\!\langle 1|.$
For the generality, let $p_{i}^{R}$ denote a noise fraction of the $i$-th measurement of a party $R\in\{A,B\}$ so that the POVM element above may be written as,
$r_{0}^{A}=\prod_{i=1}^{n}\left(1-\frac{p_{i}^{A}}{2}\right),~{}r_{1}^{A}=\prod_{i=1}^{n}\left(\frac{p_{i}^{A}}{2}\right)~{}~{}\mathrm{and}~{}~{}r_{0}^{B}=\prod_{i=1}^{m}\left(1-\frac{p_{i}^{B}}{2}\right),~{}r_{1}^{B}=\prod_{i=1}^{m}\left(\frac{p_{i}^{B}}{2}\right).$ (27)
Two parties then exploit noisy measurements purified by $(n-1)$ and $(m-1)$ additional qubits, respectively, and execute the distillation protocol. Once the protocol is successful, an unnormalized state is obtained,
$p_{succ}\rho^{\prime}$ $=F^{2}r_{\text{even}}|\phi^{+}\rangle\langle\phi^{+}|+F\left(\frac{1-F}{3}\right)\left((r_{\text{even}}+r_{\text{odd}})\Pi_{0}-r_{\text{even}}|\phi^{+}\rangle\langle\phi^{+}|\right)$
$\phantom{=}+\left(\frac{1-F}{3}\right)F\left(r_{\text{even}}(\Pi_{0}-|\phi^{+}\rangle\langle\phi^{+}|)+r_{\text{odd}}\Pi_{1}\right)+\left(\frac{1-F}{3}\right)^{2}\left(r_{\text{odd}}\Pi_{0}+(2r_{\text{even}}+r_{\text{odd}})\Pi_{1}+r_{\text{even}}|\phi^{+}\rangle\langle\phi^{+}|\right),$ (28)
where we have used $\Pi_{0}=|00\rangle\langle 00|+|11\rangle\langle 11|$ and $\Pi_{1}=|01\rangle\langle 01|+|10\rangle\langle 10|$, and the parameters are given by
$r_{\text{even}}=r_{0}^{A}r_{0}^{B}+r_{1}^{A}r_{1}^{B}~{}~{}\mathrm{and}~{}~{}r_{\text{odd}}=r_{0}^{A}r_{1}^{B}+r_{1}^{A}r_{0}^{B},$ (29)
with
$r_{\text{even}}^{(n,m)}$ $=\prod_{i=1}^{n}\left(1-\frac{p^{A}_{i}}{2}\right)\prod_{i=1}^{m}\left(1-\frac{p^{B}_{i}}{2}\right)+\prod_{i=1}^{n}\left(\frac{p^{A}_{i}}{2}\right)\prod_{i=1}^{m}\left(\frac{p^{B}_{i}}{2}\right),~{}\mathrm{and}$ (30)
$r_{\text{odd}}^{(n,m)}$ $=\prod_{i=1}^{n}\left(1-\frac{p^{A}_{i}}{2}\right)\prod_{i=1}^{m}\left(\frac{p^{B}_{i}}{2}\right)+\prod_{i=1}^{n}\left(\frac{p^{A}_{i}}{2}\right)\prod_{i=1}^{m}\left(1-\frac{p^{B}_{i}}{2}\right).$ (31)
Appendix B Entanglement distillation with noisy purification protocols: mixed states
In this section, we detail the analysis on the two-way entanglement distillation protocol with the purification of noisy measurements in which CNOT gates are noisy. We show a lower bound of the distillability:
  • i) $F>1/2$ for a noiseless case,
  • ii) $F>L^{(n,m)}$ in Eq. (17) where $L^{(n,m)}$ converges to $1/2$ when $n$ and $m$ tend to be large, and
  • iii) $F>L_{\epsilon}^{(n,m)}$, with $\epsilon$ a noise fraction in a CNOT gate, see Eq. (19), where $L_{\epsilon}^{(n,m)}$ does not converge to $1/2$ for large $n$ and $m$.
To show the item iii), it suffices to consider instances $n=m$. A lower bound found from Eq. (17)
$L_{\epsilon}^{(n)}$ $=$ $\frac{r_{even}^{(n)}+r_{odd}^{(n)}}{2(r_{even}^{(n)}-r_{odd}^{(n)})}=\frac{(r_{0}^{(n)}+r_{1}^{(n)})^{2}}{2(r_{0}^{(n)}-r_{1}^{(n)})^{2}}.$
where $r_{j}^{(n)}$ for $j=0,1$ is a probability obtained by a purified measurement with noise CNOT gates with a noise fraction $\epsilon$. Probabilities $r_{0}^{(n)}$ and $r_{1}^{(n)}$ are obtained by a recurrence relation:
$r_{0}^{(n+1)}$ $=(1-\epsilon)r_{0}^{(n)}(1-\frac{p}{2})+\frac{\epsilon}{4}(r_{0}^{(n)}+r_{1}^{(n)})~{}\mathrm{and}$
$r_{1}^{(n+1)}$ $=(1-\epsilon)r_{1}^{(n)}(\frac{p}{2})+\frac{\epsilon}{4}(r_{0}^{(n)}+r_{1}^{(n)}).$
From the relation above, we write by $s$ the limit point,
$s=\lim_{n\rightarrow\infty}\frac{r_{1}^{(n)}}{r_{0}^{(n)}},$
for which it holds that
$s=\frac{(1-\epsilon)s(\frac{p}{2})+\frac{\epsilon}{4}(1+s)}{(1-\epsilon)(1-\frac{p}{2})+\frac{\epsilon}{4}(1+s)}.$
Hence, the limit can be found as
$s$ $=$ $2(1-p)(1-\frac{1}{\epsilon})+\sqrt{5-4p(2-p)+\frac{4(1-p)^{2}}{\epsilon}(\frac{1}{\epsilon}-2)}.$
Consequently, when $n$ tends to be large, the lower bound converges as follows,
$\lim_{n\rightarrow\infty}L_{\epsilon}^{(n)}=\frac{2(1-p)^{2}(1-2\epsilon)+\epsilon^{2}(3-2(2-p)p)+\epsilon\sqrt{4(1-p)^{2}(1-2\epsilon)+\epsilon^{2}(5+4(-2+p)p)}}{4(1-\epsilon)^{2}(1-p)^{2}}$
which is strictly away from $1/2$.
Appendix C Entanglement distillation with noisy purification protocols: pure states
Suppose that we realize the purification of a measurement with $(n-1)$ additional qubits to distill entanglement from a single-copy pure state, where a POVM element is denoted by $\widetilde{Q}_{0}^{(n)}=r_{0}^{(n)}M_{0}+r_{0}^{(n)}M_{1}$. Let $p_{succ}$ denote the probability that the purification of a measurement is successful: we have
$p_{succ}~{}\rho_{n}^{(AB)}$ $=$ $\mathrm{tr}_{E}[U^{(BE)}|\psi\rangle\!\langle\psi|^{(AB)}\otimes|0\rangle\!\langle 0|^{(E)}U^{(BE)\dagger}\widetilde{Q}_{0}^{(E)}]$
$=$ $r_{0}^{(n)}2\sin^{2}\theta|\phi^{+}\rangle\!\langle\phi^{+}|+r_{1}^{(n)}(1-2\sin^{2}\theta)|11\rangle\!\langle 11|.$
In this case, we compute the fidelity as follows,
$F_{n}=\langle\phi^{+}|\rho_{n}^{(AB)}|\phi^{+}\rangle=\frac{1}{p_{succ}}\left(r_{0}^{(n)}2\sin^{2}\theta+\frac{1}{2}r_{1}^{(n)}(1-2\sin^{2}\theta)\right)~{}~{}\mathrm{where}~{}~{}p_{succ}$ $=r_{0}^{(n)}2\sin^{2}\theta+r_{1}^{(n)}(1-2\sin^{2}\theta).$
If noise fractions $p$ are identical, and CNOT gates in measurement purification have an error $\epsilon$, $r_{0}^{(n)}$ and $r_{1}^{(n)}$ can be calculated by recursive relation:
$r_{0}^{(n+1)}$ $=(1-\epsilon)r_{0}^{(n)}(1-\frac{p}{2})+\frac{\epsilon}{4}(r_{0}^{(n)}+r_{1}^{(n)})~{}\mathrm{and}$
$r_{1}^{(n+1)}$ $=(1-\epsilon)r_{1}^{(n)}(\frac{p}{2})+\frac{\epsilon}{4}(r_{0}^{(n)}+r_{1}^{(n)}),$
with a starting condition $r_{0}^{(1)}=(1-p/2)$ and $r_{1}^{(1)}=p/2$.
For $\epsilon=0$, we have $r_{0}^{(n)}=(1-\frac{p}{2})^{n}$ and $r_{1}^{(n)}=(\frac{p}{2})^{n}$. One can find that $F_{n}$ converges to $1$ as $n$ tends to be large in this case. For $\epsilon>0$, we have
$s=\lim_{n\rightarrow\infty}\frac{r_{1}^{(n)}}{r_{0}^{(n)}}=2(1-p)(1-\frac{1}{\epsilon})+\sqrt{5-4p(2-p)+\frac{4(1-p)^{2}}{\epsilon}(\frac{1}{\epsilon}-2)}$
and
$\lim_{n\rightarrow\infty}F_{n}$ $=\frac{{2\sin^{2}\theta+\frac{1}{2}s(1-2\sin^{2}\theta)}}{2\sin^{2}\theta+s(1-2\sin^{2}\theta)}$
$=\frac{2\epsilon+\left(-2+2p-2\epsilon p+\sqrt{\epsilon^{2}(5+4(-2+p)p)+4(1-p)^{2}(1-2\epsilon)}\right)\cos(2\theta)}{2\epsilon+2\left(-2+\epsilon+2p-2\epsilon p+\sqrt{\epsilon^{2}(5+4(-2+p)p)+4(1-p)^{2}(1-2\epsilon)}\right)\cos(2\theta)}.$